# Deep Generative Models in Condition and Structural Health Monitoring: Opportunities, Limitations and Future Outlook

### v2

@@ 3.1.2 DGMs for data generation
Signal processing Conventional ML/DNNs DGMs Data transformation-based/Least-squares SMOTE/CNN/RNN/LSTM DAR/VAE/GAN/Diffusion Data Augmentation Refs: Data transformation-based [62, 63, 64, 65] Refs: SMOTE-based [66, 67, 68] Refs: DAR-based [69, 70, 71, 72], VAEs [73, 74], GANs [75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90], Diffusion-based [91, 92, 93, 60, 94, 95, 96, 97, 98, 99, 100, 101, 102] Quantitative comparison: – [103] DDPM-based method surpasses GAN-based and autoencoder model in augmentation task with metric: MMD . – [104] Different GAN-based variants for wavelet transformed image augmentation: MMD, EMD, , SSIM. – [105] Comparison of LSTM, VAE, GAN-based data augmentation: Improved diagnosis performance ACC (confusion matrix). – [106] SMOTE and GAN-based data augmentation performance comparison: classification accuracy (F1-score, confusion matrix). Data Imputation Refs: Least-squares [107], Bayesian [108] Refs: KNN [109], Principal component analysis (PCA) [110], RNN [111] Refs: DAR-based [112, 113], VAEs [114], GANs [115, 116, 117, 118], Diffusion-based [119] Quantitative comparison: – [120] The proposed LSTM-based methods outperform random forest (RF), light gradient boosting machine, and multiple linear regression for dams missing data imputation task: , MAE , RMSE . – [121] The proposed Bayesian matrix learning approach outperforms PCA in SHM data imputation: higher ACC (0.95). – [122] Comparison among KNN, RF, matrix factorization, RNN, their proposed model: MAE (7.77), RMSE (10.92). – [123] Dam missing value imputation among GRU, LSTM, and their proposed model: MAE (0.593), MAPE (0.076), RMSE (0.731). – [124] GAN surpasses SVR, radial basis function network, and RNN in data imputation: lower MSE (0.0237), higher (0.9946). – [125] The proposed GAN-based method surpasses LSTM and SVR in bridge data imputation: ACC (0.9), lower execution time. – [126] The method outperforms the VAE and other GAN variants with: lower RMSE (), MAE () values, and higher . Domain Adaptation Refs: Not applicable Refs: Gaussian mixture model [127], CNN-based models [128] Refs: GANs [129, 130, 131, 132, 133, 134, 135, 136, 87], Diffusion-based [137] Quantitative comparison: – [138] Comparison among different DA methods on four SHM scenarios: ACC , F1-score (1.000). – [139] Improved diagnostic performance of proposed dynamic collaborative adversarial domain adaptation network in unsupervised DA task: recall , precision , accuracy , F1-score (classification metrics). – [140] Multi-source DA comparison among diffusion and DNN models: improved classification performance (confusion matrix, accuracy). Domain Generalization Refs: Not applicable Refs: Mete-learning techniques [141] Refs: GANs [90], Diffusion-based model [142] Quantitative comparison: – [143] The proposed adversarial domain generalization framework that outperforms LSTM, CNN-based, temporal CNN, and multiscale CNN with evaluation metrics on unknown target domain: RMSE (0.1376), MAE (0.1117). – [144] Meta-learning-based DG model outperforms domain adversarial neural network, correlation alignment, meta-learning DG, aggregation with improved metrics: accuracy (), enhanced classification performance with confusion matrix. – [145] The proposed similarity-based meta-representation learning method for DG task outperforms several baseline methods including neural network, hierarchical neural network, model-agnostic meta-learning (MAML), and hierarchical MAML with improved metric: ACC .
@@ 3.1.3 DGMs for domain adaptation and generalization
Transfer learning (TL) aims to leverage knowledge from a pre-trained model on a source task to improve performance on a different yet related target domain or task [146, 147]. Within this framework, domain adaptation (DA) and domain generalization (DG) emerge as two pivotal subfields. DA focuses on minimizing data distribution discrepancies between the source and target domains [148, 149], while DG seeks to transfer knowledge from multiple source domains to an unseen target domain without utilizing any data from the target domain. Hence, the main difference between DA and DG lies in data accessibility during training: DA can utilize target domain data during training (even when such data are scarce or limited), while DG relies solely on source domain knowledge. This makes DG particularly suitable for evaluating the robustness of the model in completely unseen working scenarios where the target domain data remain inaccessible [82]. Fig. 5 illustrates the main difference between DA and DG concepts.
Following the framework established in the previous subsection, Table 4 outlines the applications of DGMs for DA and DG within the TL framework. Apparently, signal processing methods are not applicable to either DA or DG, as they belong to transfer learning, which is indeed one category of ML methods. Nevertheless, recent work mainly focuses on GAN-based models for DA, including: (i) undamaged-to-damaged data transition under scarce damage data [129, 130, 131]; (ii) adaptation between simulation and experimental data [135, 132, 139], and (iii) adaptation among different rotating machine types, speeds, or operating conditions [133, 134, 87]).
A smaller proportion of studies [90, 142, 145, 144, 143] address DG under unknown or unseen working conditions. Meta-learning [145, 144] stands out as a complementary method to DG by training a model to adapt quickly to new tasks or environments by leveraging knowledge gained from multiple related tasks. Since DG is practical in industrial settings, where target domain data is unavailable, there is an emerging trend toward the use of diffusion model-based DG for CM/SHM systems [103]. It is noted that certain citations span multiple application categories. For instance, [90, 87, 103] proposed methodologies that combine data augmentation with domain generalization to handle unknown target domains. As shown in Table 4, these approaches inherently span both categories since they both involve augmenting scarce data while achieving data generalization to unseen domains.
@@ 3.1.4 DGMs for multimodal data fusion
In the context of multimodal data fusion in condition monitoring, the application of DGMs remains in its early stages. Notably, diffusion-based and LLM-based DGMs exhibit unique promise largely because of their Transformer-based architectures, which inherently enable the dynamic fusion of heterogeneous data types, including text, visual images, and audio inputs (known as multimodal data), into a unified latent space. The attention mechanism [35] within these Transformer frameworks serves as a foundation for cross-modal fusion. By dynamically computing attention scores across modalities, this mechanism enables context-aware weighting of input features, allowing the model to prioritize statistically salient or semantically correlated signals during fusion. These capabilities are particularly important for diffusion-based and LLM-based DGMs, as the attention mechanism not only enhances representational power but also ensures the coherence of model outputs with the joint distribution of multimodal inputs.
Fig. 6(a) provides an illustrative scheme of the multi-head attention mechanism in the Transformer network, which is a fundamental component in both Diffusion-based (e.g., Stable Diffusion [150]) and LLM-based models. Fig. 6(b) illustrates the implementation of cross-modal fusion in LLMs, where paired image and text inputs are jointly processed. In this framework, the image (or visual) inputs are sent to a feature extraction module via an image encoder to obtain latent representations, which are subsequently projected into key () and query () matrices. Meanwhile, the text caption is encoded by a text encoder (e.g., CLIP text encoder [151]) into a value () matrix, as shown in Fig. 6(a). In this way, the LLM computes attention scores by enabling the dynamic fusion of image and text modalities. Fig. 6(c) depicts the main flow of multimodal data fusion based on DGM. Various data formats, such as discrete data, continuous time series, or images, can be obtained from multimodal heterogeneous sensors (such as vibration accelerometers, stereo cameras, and acoustic emission sensors). The DGMs can operate on and dynamically fuse this information to facilitate downstream FD&AD tasks. Table 5 concludes the DGM application for multimodal data fusion.
@@
Signal processing Conventional ML/DNNs DGMs Kalman filtering/Least-squares CNN/RNN/LSTM/Transformer Diffusion/LLM Multimodal Data Fusion Refs: Kalman filtering [152, 153, 154] Refs: LSTM-based [155], Transformer [156, 157] Refs: Diffusion-based [158], LLM-based [159, 14, 160, 161, 162] Quantitative comparison: – [163] A multi-scale Transformer network was proposed with acceleration feature fusion achieving improved performance comparing to CNNs and sigle-scale Transformer network: Accuracy (96.12%), confusion matrix. – [164] The proposed Swin-Transformer-based model achieves highest diagnostic accuracy compared to ResNet18, EfficienceV2, and Vgg16 networks by fusing infrared thermal and visible light images: Accuracy , confusion matrix. Fault Diagnosis &Anomaly Detection Refs: Not applicable Refs: CNN-based [165, 166, 167], Transformer [168, 169, 156, 157, 170, 171] Refs: Diffusion-based, LLM-based [14, 160, 161, 162] Quantitative comparison: – [156] A two-stage encoder Transformer was proposed with 50% improvement in accuracy comparing to extreme gradient boosting. – [157] The temporal fusion Transformer network achieves higher accuracy compared to the autoregressive model: MAE, MSE, MAPE. – [172] The proposed model outperforms ShuffleNet, ResNet, DenseNet models with improved: average accuracy ().
@@ 3.1.5 DGMs for fault diagnosis and anomaly detection
For downstream FD&AD tasks, it needs to be noted that the CM/SHM process collects sensor data from different positions and channels with varying types and sampling rates. FD and AD were carried out based on the collected data. LLM-based diagnosis and prognostic methods have recently been developed, thanks to the development of game-changing products, such as ChatGPT, which are developed on Transformer-based architecture with the ability to create text, images, and even videos.
The basic idea behind applying LLM-based models in FD&AD tasks is that the Transformer architecture in LLMs integrates a large number of model parameters that are pre-trained on large text-image datasets. Then the pre-trained LLMs can be adapted to bearing or structural FD&AD tasks by fine-tuning the customized dataset. As shown in Fig. 7, the authors conducted fault diagnosis by using Low-Rank Adaptation (LoRA) fine-tuned GPT-2 model based on bearing vibration datasets. In Table 5, we summarize the LLM-based FD&AD tasks presented for the wind turbine system [14], and the bearing/gearbox [160, 161, 162]. Most of the publications for LLM-based FD&AD tasks focused on bearing/gearbox and were also based on LoRA fine-tuning techniques. We also detailed the different types of fine-tuning techniques in the following Section 4.1.1.
It is noted that the fine-tuning efficiency and model size of LLMs largely determine their final performance in FD&AD tasks. Zheng et al. [173] compared the diagnostic accuracy of Llama-2 models of different sizes (7B vs 13B), and reported that the 7B model can achieve satisfactory results, while larger models are generally preferred for tasks requiring higher precision or involving more complex datasets. Additional studies discussing the impact of fine-tuning strategies and model scale on LLM performance for FD&AD can be found in Refs. [174, 175, 176, 177].
@@ 3.2 Applications in industrial systems
In the previous subsection we summarize the main tasks performed by DGMs. To illustrate how these tasks transfer into practical use, this section describes DGM applications across subsystems and components in four major industrial sectors: rotating machinery, aircraft structures, wind turbines, and civil infrastructure. Fig. 8 maps base DGM families and their variants to the tasks they perform and the subsystems they enable in CM/SHM. From the figure, it can be observed that GAN and its variants dominate in data generation task and most of tasks are applied to rotating machinery and civil infrastructure sectors. The diffusion-based models and LLMs remain in the early stages of industrial practice.
@@ 3.2.1 Rotating machinery
Rotating machinery is a major and critical component of many mechanical systems, serving primarily to generate, transmit or regulate rotational motion and torque [178]. A representative example is the gearbox, a mechanical device designed to adjust the rotational speed and output power to the desired level. Inside the gearbox, gears transmit torque through precise meshing, while bearings support power transmission by reducing friction and stabilizing the rotating shafts. These components, however, are highly susceptible to unexpected failures: bearings are likely to experience inner race, outer race, or rolling element faults [78], whereas gears are prone to surface fatigue (pitting, case crushing, and spalling) and even tooth breakage [179]. Therefore, continuous monitoring is essential to detect problems and predict potential failures.
As shown in Table 6, DGMs have been extensively applied to vibration acceleration signals, supporting a range of tasks discussed in Section 3.1. At the component or subsystem level, gearboxes and bearings are the most representative components in rotating machinery. GAN and diffusion-based models are commonly employed for data generation, domain adaptation, and generalization tasks. Recently, LLM-based approaches for multimodal fusion and fault diagnosis have emerged as a growing trend. The text prompting functionality enabled by LLMs allows the recording and interpretation of metadata, such as operational dialogues and state changes, which other DGMs cannot process, thereby offering new opportunities for knowledge-enhanced condition monitoring.

### v3

@@ 3.1.2 DGMs for data generation
Signal processing Conventional ML/DNNs DGMs Data transformation-based/Least-squares SMOTE/CNN/RNN/LSTM DAR/VAE/GAN/Diffusion Data Augmentation Refs: Data transformation-based [9259076, NGUYEN2022114172, 9673114, OH202072] Refs: SMOTE-based [JANSSEN2020115483, 8815712, 8737894] Refs: DAR-based [CHEN2023116063, ZHANG2022109175, djemili2024wind, Chen10081265], VAEs [LI2025110027, li2023machine], GANs [LI2025112062, Xiaoming2021, 10058512, WANG2025110760, ZHUANG2024111186, WANG2023106872, LIU2022110888, WAN2025109614, LI2022101552, LI20242958, ZHANG2024114795, SHEN2024109159, Sun10504387, WEN2024111663, XU2023107063, GUO2025110854], Diffusion-based [ZHENG2025119685, SHEN2024109299, DENG2025116595, SHU2025111783, 10750268, MUELLER2024107696, ZHAO2025109520, YU2024110343, Wang0309714, HerTerng, YANG2025110312, YI2024111481, WANG2024110394] Quantitative comparison: – [GUO2025115951] DDPM-based method surpasses GAN-based and autoencoder model in augmentation task with metric: MMD . – [Fu_2023] Different GAN-based variants for wavelet transformed image augmentation: MMD, EMD, , SSIM. – [10136604] Comparison of LSTM, VAE, GAN-based data augmentation: Improved diagnosis performance ACC (confusion matrix). – [10239215] SMOTE and GAN-based data augmentation performance comparison: classification accuracy (F1-score, confusion matrix). Data Imputation Refs: Least-squares [van2015time], Bayesian [QiAngWang2022] Refs: KNN [chirici2016meta], Principal component analysis (PCA) [9851940], RNN [YE2024] Refs: DAR-based [ZHANG2022108718, WANG2023110703], VAEs [Ahang_2024], GANs [QU2020106610, ZHENG2025119694, GAO2023105277, Huachen2022], Diffusion-based [NEURIPS2024_cb1ba6a4] Quantitative comparison: – [LI2021109377] The proposed LSTM-based methods outperform random forest (RF), light gradient boosting machine, and multiple linear regression for dams missing data imputation task: , MAE , RMSE . – [REN2021107734] The proposed Bayesian matrix learning approach outperforms PCA in SHM data imputation: higher ACC (0.95). – [tan2024missing] Comparison among KNN, RF, matrix factorization, RNN, their proposed model: MAE (7.77), RMSE (10.92). – [song2024missing] Dam missing value imputation among GRU, LSTM, and their proposed model: MAE (0.593), MAPE (0.076), RMSE (0.731). – [HOU2022111206] GAN surpasses SVR, radial basis function network, and RNN in data imputation: lower MSE (0.0237), higher (0.9946). – [GAO2022112095] The proposed GAN-based method surpasses LSTM and SVR in bridge data imputation: ACC (0.9), lower execution time. – [WANG2025110973] The method outperforms the VAE and other GAN variants with: lower RMSE (), MAE () values, and higher . Domain Adaptation Refs: Not applicable Refs: Gaussian mixture model [gardner2022domain], CNN-based models [chen2023deep] Refs: GANs [SOLEIMANIBABAKAMALI2023110404, LULECI2023110370, LULECI2023106146, Lou9789138, ir202307, WANG2023455, GE2024111236, CACERESCASTELLANOS20237746, Sun10504387], Diffusion-based [ZHANG2025107031] Quantitative comparison: – [GARDNER2020106550] Comparison among different DA methods on four SHM scenarios: ACC , F1-score (1.000). – [WANG2025110662] Improved diagnostic performance of proposed dynamic collaborative adversarial domain adaptation network in unsupervised DA task: recall , precision , accuracy , F1-score (classification metrics). – [XU2024112396] Multi-source DA comparison among diffusion and DNN models: improved classification performance (confusion matrix, accuracy). Domain Generalization Refs: Not applicable Refs: Mete-learning techniques [finn2017modelagnostiCMetalearningfastadaptation] Refs: GANs [GUO2025110854], Diffusion-based model [LIU2025116989] Quantitative comparison: – [Zhuang9759507] The proposed adversarial domain generalization framework that outperforms LSTM, CNN-based, temporal CNN, and multiscale CNN with evaluation metrics on unknown target domain: RMSE (0.1376), MAE (0.1117). – [10091197] Meta-learning-based DG model outperforms domain adversarial neural network, correlation alignment, meta-learning DG, aggregation with improved metrics: accuracy (), enhanced classification performance with confusion matrix. – [10688397] The proposed similarity-based meta-representation learning method for DG task outperforms several baseline methods including neural network, hierarchical neural network, model-agnostic meta-learning (MAML), and hierarchical MAML with improved metric: ACC .
@@ 3.1.3 DGMs for domain adaptation and generalization
Transfer learning (TL) aims to leverage knowledge from a pre-trained model on a source task to improve performance on a different yet related target domain or task [wang2022generalizing, YANG2024115387]. Within this framework, domain adaptation (DA) and domain generalization (DG) emerge as two pivotal subfields. DA focuses on minimizing data distribution discrepancies between the source and target domains [XU2024102650, Kouw_2021], while DG seeks to transfer knowledge from multiple source domains to an unseen target domain without utilizing any data from the target domain. Hence, the main difference between DA and DG lies in data accessibility during training: DA can utilize target domain data during training (even when such data are scarce or limited), while DG relies solely on source domain knowledge. This makes DG particularly suitable for evaluating the robustness of the model in completely unseen working scenarios where the target domain data remain inaccessible [WAN2025109614]. Fig. 5 illustrates the main difference between DA and DG concepts.
Following the framework established in the previous subsection, Table 4 outlines the applications of DGMs for DA and DG within the TL framework. Apparently, signal processing methods are not applicable to either DA or DG, as they belong to transfer learning, which is indeed one category of ML methods. Nevertheless, recent work mainly focuses on GAN-based models for DA, including: (i) undamaged-to-damaged data transition under scarce damage data [SOLEIMANIBABAKAMALI2023110404, LULECI2023110370, LULECI2023106146]; (ii) adaptation between simulation and experimental data [GE2024111236, Lou9789138, WANG2025110662], and (iii) adaptation among different rotating machine types, speeds, or operating conditions [ir202307, WANG2023455, Sun10504387]).
A smaller proportion of studies [GUO2025110854, LIU2025116989, 10688397, 10091197, Zhuang9759507] address DG under unknown or unseen working conditions. Meta-learning [10688397, 10091197] stands out as a complementary method to DG by training a model to adapt quickly to new tasks or environments by leveraging knowledge gained from multiple related tasks. This is particularly relevant because many DGM-based DG approaches rely primarily on statistical learning from available data, without explicitly incorporating physical knowledge of the system. Since DG is highly desirable in industrial settings, where target domain data is unavailable, there is an emerging trend toward the use of diffusion model-based DG for CM/SHM systems [GUO2025115951]. It is noted that certain citations span multiple application categories. For instance, [GUO2025110854, Sun10504387, GUO2025115951] proposed methodologies that combine data augmentation with domain generalization to handle unknown target domains. As shown in Table 4, these approaches inherently span both categories since they both involve augmenting scarce data while achieving data generalization to unseen domains.
@@ 3.1.4 DGMs for multimodal data fusion
In the context of multimodal data fusion in condition monitoring, the application of DGMs remains in its early stages. Notably, diffusion-based and LLM-based DGMs exhibit unique promise largely because of their Transformer-based architectures, which inherently enable the dynamic fusion of heterogeneous data types, including text, visual images, and audio inputs (known as multimodal data), into a unified latent space. The attention mechanism [vaswani2023attentionneed] within these Transformer frameworks serves as a foundation for cross-modal fusion. By dynamically computing attention scores across modalities, this mechanism enables context-aware weighting of input features, allowing the model to prioritize statistically salient or semantically correlated signals during fusion. These capabilities are particularly important for diffusion-based and LLM-based DGMs, as the attention mechanism not only enhances representational power but also ensures the coherence of model outputs with the joint distribution of multimodal inputs.
Fig. 6(a) provides an illustrative scheme of the multi-head attention mechanism in the Transformer network, which is a fundamental component in both Diffusion-based (e.g., Stable Diffusion [rombach2022highresolution]) and LLM-based models. Fig. 6(b) illustrates the implementation of cross-modal fusion in LLMs, where paired image and text inputs are jointly processed. In this framework, the image (or visual) inputs are sent to a feature extraction module via an image encoder to obtain latent representations, which are subsequently projected into key () and query () matrices. Meanwhile, the text caption is encoded by a text encoder (e.g., CLIP text encoder [radford2021learning]) into a value () matrix, as shown in Fig. 6(a). In this way, the LLM computes attention scores by enabling the dynamic fusion of image and text modalities. Fig. 6(c) depicts the main flow of multimodal data fusion based on DGM. Various data formats, such as discrete data, continuous time series, or images, can be obtained from multimodal heterogeneous sensors (such as vibration accelerometers, stereo cameras, and acoustic emission sensors). The DGMs can operate on and dynamically fuse this information to facilitate downstream FD&AD tasks. Table 5 concludes the DGM application for multimodal data fusion.
@@
Signal processing Conventional ML/DNNs DGMs Kalman filtering/Least-squares CNN/RNN/LSTM/Transformer Diffusion/LLM Multimodal Data Fusion Refs: Kalman filtering [PAL2024111482, ZHU2023116573, SHEN2023100442] Refs: LSTM-based [Jichen2025], Transformer [10922758, FALCHI2024111382] Refs: Diffusion-based [YAO2024111397], LLM-based [electronics13244912, ZHANG2025124378, TAO2025112127, PANG2024110312, qaid2024fdllmlargelanguagemodel] Quantitative comparison: – [CHEN2025111813] A multi-scale Transformer network was proposed with acceleration feature fusion achieving improved performance comparing to CNNs and sigle-scale Transformer network: Accuracy (96.12%), confusion matrix. – [YingHan2025] The proposed Swin-Transformer-based model achieves highest diagnostic accuracy compared to ResNet18, EfficienceV2, and Vgg16 networks by fusing infrared thermal and visible light images: Accuracy , confusion matrix. Fault Diagnosis &Anomaly Detection Refs: Not applicable Refs: CNN-based [liu2022data, JANA2022108723, RUAN2023101877], Transformer [li2022structural, 9625031, 10922758, FALCHI2024111382, WANG2024111067, JIN2022379] Refs: Diffusion-based, LLM-based [ZHANG2025124378, TAO2025112127, PANG2024110312, qaid2024fdllmlargelanguagemodel] Quantitative comparison: – [10922758] A two-stage encoder Transformer was proposed with 50% improvement in accuracy comparing to extreme gradient boosting. – [FALCHI2024111382] The temporal fusion Transformer network achieves higher accuracy compared to the autoregressive model: MAE, MSE, MAPE. – [10103569] The proposed model outperforms ShuffleNet, ResNet, DenseNet models with improved: average accuracy ().
@@ 3.1.5 DGMs for fault diagnosis and anomaly detection
For downstream FD&AD tasks, it should be clarified that LLM-based models play a fundamentally different role compared with generative models such as VAE, GAN, or Diffusion models. In VAE- or GAN-based frameworks, DGMs are typically employed for signal-level tasks, such as data generation, feature extraction, or anomaly detection based on reconstruction error. These models primarily operate on the statistical distribution of sensor signals. In contrast, LLM-based approaches are not used as data generators for sensor-level signals. Instead, they function as pretrained representation learning and reasoning modules that are fine-tuned to adapt to diagnostic tasks. The core advantage of LLMs lies in their large-scale pretraining on diverse multimodal data, enabling them to perform contextual reasoning, cross-modal understanding, and decision support.
The basic idea behind applying LLM-based models in FD&AD tasks is that the Transformer architecture in LLMs integrates a large number of model parameters that are pre-trained on large text-image datasets. Then the pre-trained LLMs can be adapted to bearing or structural FD&AD tasks by fine-tuning the customized dataset. As shown in Fig. 7, the authors conducted fault diagnosis by using Low-Rank Adaptation (LoRA) fine-tuned GPT-2 model based on bearing vibration datasets. In Table 5, we summarize the LLM-based FD&AD tasks presented for the wind turbine system [ZHANG2025124378], and the bearing/gearbox [TAO2025112127, PANG2024110312, qaid2024fdllmlargelanguagemodel]. Most of the publications for LLM-based FD&AD tasks focused on bearing/gearbox and were also based on LoRA fine-tuning techniques. We also detailed the different types of fine-tuning techniques in the following Section LABEL:section4.1.1.
It is noted that the fine-tuning efficiency and model size of LLMs largely determine their final performance in FD&AD tasks. Zheng et al. [ZHENG2024110382] compared the diagnostic accuracy of Llama-2 models of different sizes (7B vs 13B), and reported that the 7B model can achieve satisfactory results, while larger models are generally preferred for tasks requiring higher precision or involving more complex datasets. Additional studies discussing the impact of fine-tuning strategies and model scale on LLM performance for FD&AD can be found in Refs. [chen2025large, j2024finetuningllmenterprise, 10774330, LIN2025103208].
@@ 3.2 Applications in industrial systems
In the previous subsection, we summarize the main tasks performed by DGMs. To illustrate how these tasks transfer into practical use, this section describes DGM applications across subsystems and components in four major industrial sectors: rotating machinery, aircraft structures, wind turbines, and civil infrastructure. Fig. 8 maps base DGM families and their variants to the tasks they perform and the subsystems they enable in CM/SHM. From the figure, it can be observed that GAN and its variants dominate in data generation task and most of tasks are applied to rotating machinery and civil infrastructure sectors. The diffusion-based models and LLMs remain in the early stages of industrial practice, as they require substantial computational resources and large datasets.
@@ 3.2.1 Rotating machinery
Rotating machinery constitutes a fundamental and critical component of numerous mechanical systems, primarily serving to generate, transmit, or regulate rotational motion and torque [LEI2013108]. A representative example is the gearbox, which is designed to adjust rotational speed and output power to specified levels. Within the gearbox, gears transmit torque through precise meshing, while bearings support power transmission by reducing friction and stabilizing the rotating shafts. These components, however, are highly susceptible to unexpected failures: bearings are likely to experience inner race, outer race, or rolling element faults [WANG2025110760], whereas gears are prone to surface fatigue (pitting, case crushing and spalling) and even tooth breakage [SINGHRAGHAV2023104020]. Continuous monitoring is therefore essential for timely fault detection and reliable failure prediction.
As shown in Table LABEL:table_6, DGMs have been widely applied to vibration signals, supporting a variety of tasks discussed in Section 3.1. At the component or subsystem level, gearboxes and bearings are the most representative components in rotating machinery. GAN and diffusion-based models are commonly used for data generation, domain adaptation, and generalization tasks. Recently, LLM-based approaches for multimodal fusion and fault diagnosis have emerged as a growing trend. The text prompting capability enabled by LLMs allows the recording and interpretation of metadata, including operational logs and state transitions, which cannot be directly processed by other DGMs. This capability provides new opportunities for knowledge-enhanced condition monitoring.